(-2<x<3 21 Graph the feasible region for the system-15y 35 (2x + y<6
Answer #1

We have to find the feasible region.

For this we plot the all the lines and find the common region according to the inequalities.

The feasible region is the shaded region excluding the line

x= - 2 and 2x + y =6

1 0 Given System 5-2<x<3 | 21+ < - 10 ) з 2х+4= 1, 21+}= 6 чер л-ахів х=-2 (0,6 (-2,5 TIMESLIFEI.CLASSIFIED AND NOIDA TIMES C